OpenStudy (anonymous):

How would you determine the percentage saved if item A cost $15,500 and item B cost $24,300?

OpenStudy (anonymous):

Cross multiplying we have: 15500/24300=x/100, x = 63% => percentage saved is 100%-63% = 47%